OpenStudy (anonymous):

how to find a reference angle for Θ=-155 degrees? i have a test tomorrow and would prefer a walk through instead of a direct answer

OpenStudy (mertsj):

Which quadrant is the angle in?

OpenStudy (anonymous):

Two or more coterminal angles have the same reference angle. Assume angle A is postive and less than 360 o (2Pi), we have 4 possible cases: 1. If angle A is in quadrant I then the reference angle A r = A. 2. If angle A is in quadrant II then the reference angle A r = 180 o - A if A is given degrees and A r = Pi - A if A is given in radians. 3. If angle A is in quadrant III then the reference angle A r = A - 180 o if A is given degrees and A r = A - Pi if A is given in radians. 4. If angle A is in quadrant IV then the reference angle A r = 360 o - A if A is given degrees and A r = 2Pi - A if A is given in radians.

OpenStudy (anonymous):

**EXAMPLE** Find the reference angle to angle A = 120 o. Solution: Angle A is in quadrant II and the reference angle is given by A r = 180o - 120o = 60o

OpenStudy (anonymous):

so, i add 360 to find the coterminal angle, which is 205 and in quadrant 3. so tinge it is quadrant 3 i take 205-180=25o? is that right?

OpenStudy (mertsj):

Here is the simple way. Graph the angle and figure out how many degrees it is from the x axis. That is the reference angle.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

99